# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v11

@@ 5. Main Result and Proof
Now, although the right-hand side of   5.4 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
@@
$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
Separating the bidegrees in   5.5 yields the two independent conditions
@@
Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a non‑empty open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants). Denote this open set by $\mathcal{U}_{2}\subset M_{2}$.
Now fix any point $y_{0}\in\mathcal{U}_{2}$. For this fixed $y_{0}$, the first equation in (5.6) becomes:
@@
Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ where they are pointwise linearly independent.
Fix $x_{0}\in\mathcal{U}_{1}$. The second equation in (5.6) then gives for all $y\in M_{2}$:
@@
Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$). The lemma guarantees a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
@@
An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}$ such that
@@
Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$. By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent. Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a non‑empty open set $\mathcal{U}_{2}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a non‑empty open subset of $M=M_{1}\times M_{2}$. For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
@@
$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
linearly independent curvature operators. By the Ambrose–Singer theorem [2], the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C})$ is generated by curvature values. The off-diagonal component $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ (Definition 3.2) contains all off-diagonal projections of elements of $\mathfrak{hol}_{p}(\nabla^{C})$. Since the $r^{\sharp}$ operators constructed above are off-diagonal curvature operators, they belong to $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$, and therefore
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
In particular, if $r^{\sharp}>0$, the holonomy representation is irreducible with respect to $V_{1}\oplus V_{2}$.
@@
The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata. ∎
@@
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
@@
In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
@@
In Lemma 5.1 the reduction to an analytic metric allows us to apply the identity principle for analytic functions: if an analytic section of a vector bundle vanishes on an open set, it vanishes identically. This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms. For a smooth metric, we approximate by analytic metrics, prove the existence of a non‑empty open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.

### v12

@@ 5. Main Result and Proof
Now, although the right-hand side of   5.12 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
@@
$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
Separating the bidegrees in   5.13 yields the two independent conditions
@@
Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a dense open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants). Denote this open set by $\mathcal{U}_{2}^{2,1}\subset M_{2}$.
Now fix any point $y_{0}\in\mathcal{U}_{2}^{2,1}$. For this fixed $y_{0}$, the first equation in (5.14) becomes:
@@
Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ where they are pointwise linearly independent.
Fix $x_{0}\in\mathcal{U}_{1}^{2,1}$. The second equation in (5.14) then gives for all $y\in M_{2}$:
@@
Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$). The lemma guarantees a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}^{2,1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
@@
An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}^{2,1}$ such that
@@
Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$. By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}^{2,1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent. Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a dense open set $\mathcal{U}_{2}^{2,1}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}^{2,1}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
By a symmetric argument applied to the (1,2)-part, we obtain dense open sets for the corresponding forms on each factor. By Lemma 5.1, each of these open sets is dense (its complement has empty interior by unique continuation). Define
$\mathcal{U}_{1}:=\mathcal{U}_{1}^{(2,1)}\cap\mathcal{U}_{1}^{(1,2)}\subset M_{1},\qquad\mathcal{U}_{2}:=\mathcal{U}_{2}^{(2,1)}\cap\mathcal{U}_{2}^{(1,2)}\subset M_{2},$
where $\mathcal{U}_{1}^{(2,1)}$ and $\mathcal{U}_{1}^{(1,2)}$ are the dense open sets on $M_{1}$ arising from the $(2,1)$ and $(1,2)$ components respectively, and similarly for $M_{2}$. Since $M_{1}$ and $M_{2}$ are compact (hence complete) metric spaces, finite intersections of dense open sets are dense, hence $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are dense open subsets of $M_{1}$ and $M_{2}$ respectively.
Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a dense open subset of $M=M_{1}\times M_{2}$. For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
@@
$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}$
linearly independent curvature operators (each having non-trivial off-diagonal projection).
It remains to produce $r^{\sharp}$ linearly independent purely off-diagonal elements of the holonomy algebra. Since in the admissible setting $r^{\sharp}_{2,1}\leq 1$ and $r^{\sharp}_{1,2}\leq 1$, we have $r-\dim\mathcal{K}\leq 2$; the value $r-\dim\mathcal{K}=2$ occurs only when $r^{\sharp}_{2,1}=r^{\sharp}_{1,2}=1$. In that case, constructing two linearly independent purely off-diagonal elements at a single point would require $p\in\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, but the bidegree components $T^{2,1}$ and $T^{1,2}$ may have disjoint pointwise support, so this intersection can be empty. However, the sets $\mathcal{W}^{2,1}\cap\mathcal{U}$ and $\mathcal{W}^{1,2}\cap\mathcal{U}$ are individually non-empty open subsets of $M$ (each is a non-empty open set intersected with a dense open set), and on each one the construction below produces at least one purely off-diagonal holonomy element. This motivates the definition $r^{\sharp}:=\min(r-\dim\mathcal{K},\,1)$.
Without loss of generality assume $r^{\sharp}_{2,1}\geq r^{\sharp}_{1,2}$ (the opposite case is symmetric). Define
$\mathcal{V}\;:=\;\mathcal{W}^{2,1}\cap\,\mathcal{U},$
where $\mathcal{W}^{2,1}$ is the non-empty open set introduced in (5.11) and $\mathcal{U}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is the dense open subset constructed above. Since $\mathcal{W}^{2,1}$ is non-empty and open and $\mathcal{U}$ is dense, $\mathcal{V}$ is a non-empty open subset of $M$.
Consider the mixed curvature span
$\mathcal{S}_{p}\;:=\;\operatorname{span}\bigl\{R^{C}(X,Y):X\in V_{1},\;Y\in V_{2}\bigr\}.$
By the Ambrose–Singer theorem [3], $\mathcal{S}_{p}\subseteq\mathfrak{hol}_{p}(\nabla^{C})$. We now show that $\mathcal{S}_{p}$ contains at least one purely off-diagonal element for every $p\in\mathcal{V}$.
Suppose $r^{\sharp}_{2,1}=1$ and fix $p\in\mathcal{V}$. Since $p\in\mathcal{W}^{2,1}$, we have $h(p)\neq 0$ and $\tau_{p}\neq 0$. When $p\notin\mathcal{W}^{1,2}$ (i.e. $T^{1,2}_{p}=0$), the off-diagonal projection receives contributions from $T^{2,1}$ alone, and the argument proceeds exactly as in the single-component case below. We therefore concentrate on the more delicate situation $p\in\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, where both $T^{2,1}_{p}\neq 0$ and $T^{1,2}_{p}\neq 0$.
Write $\tilde{h}=|T^{1,2}|_{p}$, $\tilde{\sigma}=T^{1,2}_{p}/|T^{1,2}|_{p}$, and choose adapted orthonormal bases $\{e_{1},e_{2}=J_{1}e_{1}\}$ of $V_{1}$ and $\{f_{1},f_{2}=J_{2}f_{1}\}$ of $V_{2}$ with
$e_{1}=\frac{\tilde{\sigma}^{\sharp}}{|\tilde{\sigma}|},\qquad f_{1}=\frac{\tau^{\sharp}}{|\tau|},$
so that $\tau=a\,f_{1}^{*}$ and $\tilde{\sigma}=c\,e_{1}^{*}$ with $a=|\tau|>0$ and $c=|\tilde{\sigma}|>0$. In these bases the diagonal and off-diagonal projections of $R^{C}(X,Y)$ for mixed inputs $X\in V_{1}$, $Y\in V_{2}$, acting on $Z\in V_{1}$, decompose as
$R^{C}(X,Y)Z\;=\;\underbrace{\Phi(X,Y)}_{\in\,\mathfrak{g}_{0}}(Z)\;+\;\underbrace{\Xi(X,Y)}_{\in\,\mathfrak{g}_{1}}(Z),$ (5.18)
where $\mathfrak{g}_{0}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$ and $\mathfrak{g}_{1}=\operatorname{Hom}(V_{1},V_{2})\oplus\operatorname{Hom}(V_{2},V_{1})$. The diagonal projection $\Phi$ takes the form $\Phi(X,Y)=\lambda_{\mathrm{tot}}(X,Y)\,J_{1}+\mu_{\mathrm{tot}}(X,Y)\,J_{2}$, so $\ker\Phi=\ker\lambda_{\mathrm{tot}}\cap\ker\mu_{\mathrm{tot}}$ has dimension at least $2$ in $V_{1}\otimes V_{2}\cong\mathbb{R}^{4}$. Note that the cross-terms $K_{X}^{(2,1)}\circ K_{Y}^{(1,2)}$ and $K_{Y}^{(1,2)}\circ K_{X}^{(2,1)}$ map $V_{1}\to V_{1}$ (direct computation), so they contribute only to $\Phi$ and not to $\Xi$.
The off-diagonal map $\Xi(X,Y)(Z)\in V_{2}$ receives four types of contributions: a linear term and a quadratic term from each of $T^{2,1}$ and $T^{1,2}$. We decompose $\Xi=\Xi_{\mathrm{quad}}+\Xi_{\mathrm{lin}}$, where $\Xi_{\mathrm{quad}}$ collects the purely quadratic torsion terms (depending only on the pointwise values of $T$) and $\Xi_{\mathrm{lin}}$ collects the terms involving covariant derivatives of $T$. In the adapted bases with $\tau=a\,f_{1}^{*}$ and $\tilde{\sigma}=c\,e_{1}^{*}$, a direct calculation shows that the matrix of $\Xi_{\mathrm{quad}}$ with respect to the ordered basis $\{e_{1}\otimes f_{1},\,e_{1}\otimes f_{2},\,e_{2}\otimes f_{1},\,e_{2}\otimes f_{2}\}$ of $V_{1}\otimes V_{2}$ and the identification $\operatorname{Hom}(V_{1},V_{2})\cong\mathbb{R}^{4}$ is
$\Xi_{\mathrm{quad}}\;=\;\operatorname{diag}\!\Bigl(\frac{a^{2}h^{2}+c^{2}\tilde{h}^{2}}{4},\;\frac{c^{2}\tilde{h}^{2}}{4},\;\frac{a^{2}h^{2}}{4},\;0\Bigr).$ (5.19)
[…]
It follows that the three basis elements $e_{1}\otimes f_{1}$, $e_{1}\otimes f_{2}$, $e_{2}\otimes f_{1}$ are not in $\ker\Xi_{p}$ at any point of $\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, regardless of the values of the covariant derivatives of $T$.
We now produce the required purely off-diagonal element. Adding derivative-dependent terms to $\Xi_{\mathrm{quad}}$ can lower the rank of $\Xi$ from $3$ to at most $2$ (rank $1$ or 0 would force some diagonal entry of $\Xi$ to vanish, contradicting the derivative-independence just established). If $\operatorname{rank}\Xi_{p}\geq 3$, then $\dim\ker\Xi_{p}\leq 1<2\leq\dim\ker\Phi_{p}$, so $\ker\Phi_{p}\not\subset\ker\Xi_{p}$; picking any $\xi\in\ker\Phi_{p}\setminus\ker\Xi_{p}$ gives $R^{C}(\xi)\in\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\setminus\{0\}$.
[…]
To see that this constraint is not identically satisfied, observe that the condition $\operatorname{rank}\Xi_{p}\leq 2$ is equivalent to the vanishing of all $3\times 3$ minors of $\Xi_{p}$; since $\Xi_{\mathrm{quad}}$ has rank $3$, these minors evaluated at zero transverse derivatives are non-zero, hence the minors are not identically zero as functions of the transverse derivatives, and their vanishing imposes non-trivial conditions. At any point where these conditions fail—that is, where $\operatorname{rank}\Xi_{p}=3$—the dimensional argument above applies directly. Such points exist: on flat tori, harmonic $1$-forms are parallel by the Bochner theorem, so if the harmonic part of $h\tau$ dominates, the transverse derivatives $\nabla^{g_{2}}(h\tau)$ are small and the rank is $3$; on non-flat factors ($\kappa_{i}\neq 0$), a direct argument using the Gauss curvature contributions to $\mathfrak{hol}_{p}\cap\mathfrak{g}_{0}$ yields $\mathfrak{hol}_{p}\supset\mathfrak{g}_{0}$, and any element of $\mathcal{S}_{p}$ with $\pi_{1}\neq 0$ then produces a purely off-diagonal element by subtraction. In all cases, the set $\{p\in\mathcal{W}^{2,1}\cap\mathcal{U}:\operatorname{rank}\Xi_{p}\geq 3\}$ is open (the rank of a smooth linear map is lower semicontinuous) and dense. Density is established by the same Greene–Jacobowitz approximation scheme used in Lemma 5.1: for a real-analytic metric, the locus $\{\operatorname{rank}\Xi\leq 2\}$ is a proper analytic subset (since $\operatorname{rank}\Xi_{\mathrm{quad}}=3$ furnishes a point in the jet space where all relevant minors are nonzero, and the identity principle for analytic functions applies); for a smooth metric, $C^{2}$-approximation by analytic metrics transfers the density to the limit. On this dense open set, $\dim\ker\Xi_{p}\leq 1<2\leq\dim\ker\Phi_{p}$, so $\ker\Phi_{p}\not\subset\ker\Xi_{p}$.
In either case, we obtain a nonzero $\xi\in\ker\Phi_{p}\setminus\ker\Xi_{p}$ and hence
$R^{C}(\xi)\;\in\;\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\setminus\{0\}.$ (5.20)
When $p\notin\mathcal{W}^{1,2}$ (so $T^{1,2}_{p}=0$), the diagonal projection reduces to $\lambda(X,Y)\,J_{1}$ and the off-diagonal map receives contributions only from $T^{2,1}$. The original argument applies: pick $Y\in V_{2}$ with $\tau(Y)\neq 0$ and choose $X_{0}\in\ker(X\mapsto\lambda(X,Y))$; then $R^{C}(X_{0},Y)$ is purely off-diagonal and nonzero by the orthogonality of $g(J_{1}X_{0},\cdot)$ and $g(X_{0},\cdot)$. The symmetric argument, exchanging the roles of $V_{1}$ and $V_{2}$, applies when $r^{\sharp}_{1,2}=1$ and produces a nonzero purely off-diagonal element in $\operatorname{Hom}(V_{2},V_{1})$. Each construction independently yields at least one purely off-diagonal element on its respective open set.
Since $\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\subseteq\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)=\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ (Definition 3.2), we conclude that for every $p\in\mathcal{V}$,
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;=\;\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr).$ (5.21)
@@
In particular, if $r^{\sharp}>0$, the holonomy representation at every point of the non-empty open set $\mathcal{V}$ is irreducible with respect to $V_{1}\oplus V_{2}$, therefore, we have global irreducibility. The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}^{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata. ∎
@@
Note that the following example lies outside the admissible class but illustrates why the non-parallel torsion condition is essential.
@@
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr),$
and irreducibility is forced precisely when $r^{\sharp}>0$, i.e. when
@@
In other words, a non‑zero value of the corrected rank $r^{\sharp}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
@@
In Lemma 5.1 the reduction to an analytic metric allows us to apply the identity principle for analytic functions: if an analytic section of a vector bundle vanishes on an open set, it vanishes identically. This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms. For a smooth metric, we approximate by analytic metrics, prove the existence of a dense open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a dense open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.
@@
Remark 5.7 (Dichotomy of the obstruction).
For admissible product manifolds, the obstruction space $\mathcal{K}$ is controlled by a single topological datum: the vanishing or non-vanishing of $\mathcal{P}_{1}(M_{i})$. Since $\mathcal{P}_{2}(M_{i})=\mathbb{R}\cdot\operatorname{vol}_{M_{i}}$ for every compact oriented surface, the obstruction kernels reduce to
$\ker\Psi_{p}\neq\{0\}\;\Longleftrightarrow\;\beta\in\mathcal{P}_{1}(M_{2}),\qquad\ker\widetilde{\Psi}_{p}\neq\{0\}\;\Longleftrightarrow\;\alpha\in\mathcal{P}_{1}(M_{1}),$
where $\alpha$, $\beta$ are the harmonic forms appearing in the minimal decomposition of $\omega]_{\mathrm{mixed}}$. A compact oriented surface $F$ satisfies $\mathcal{P}_{1}(F)=\{0\}$ if and only if either $g(F)\geq 2$ (by the Poincaré–Hopf theorem: $\chi(F)\neq 0$ forbids nowhere-vanishing vector fields, hence parallel $1$-forms) or $g(F)=1$ and the metric is not flat (a non-zero parallel $1$-form on a surface forces the Gaussian curvature to vanish identically).
This yields a sharp dichotomy. When at least one factor has genus $g\geq 2$, the obstruction space $\mathcal{K}$ is at most one-dimensional regardless of the metric, and $r^{\sharp}=1$ whenever the mixed class $\omega]_{\mathrm{mixed}}$ is non-trivial in the corresponding Künneth component; in particular, on $\Sigma_{g_{1}}\times\Sigma_{g_{2}}$ with $g_{1},g_{2}\geq 2$, the bound $r^{\sharp}=1$ is purely topological and forces irreducibility for every cohomologically calibrated connection with non-parallel torsion representing a non-trivial mixed class, independently of the choice of product metric. The only setting where the metric controls the transition $r^{\sharp}=0\leftrightarrow r^{\sharp}=1$ is when the relevant factor is a torus $T^{2}$: the parallel-form stratum $\mathcal{P}_{1}(T^{2},g)$ jumps from $\mathbb{R}^{2}$ to $\{0\}$ as soon as the Gaussian curvature becomes non-identically-zero, and this jump is the sole mechanism by which the bound changes.
Finally, we note that the untruncated quantity $r-\dim\mathcal{K}$ counts the number of independent mixed directions that are not absorbed by parallel forms. In the admissible setting this integer can reach the value $2$ (when both Künneth components contribute and $\mathcal{K}=\{0\}$), but Theorem 5.2 transfers only $\min(r-\dim\mathcal{K},\,1)$ of these directions into $\mathfrak{hol}^{\mathrm{off}}$ as purely off-diagonal elements. The truncation arises because the current proof constructs at most one purely off-diagonal holonomy element per bidegree component at a single point; the question of whether this truncation can be removed is discussed in Section 7.